# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

### v2

@@ 3.5 Evaluation metrics
We introduce instance-level metrics extending traditional REC metrics to address multi-target challenges. Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples. However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to hundreds of instances per expression in our task. To address this limitation, these new metrics provide a more granular evaluation of a model’s capability to identify individual target objects.
Instance-level metrics: Accinst. and F1inst. provide a fine-grained evaluation of the ability to precisely localize individual target objects. We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes. An IoU $\geq$ 0.5 is considered a true positive (TP), otherwise it is a false positive (FP). Unmatched true bounding boxes are counted as false negatives (FN). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate:
@@
[…]
Image-level metrics: Accimg. and F1img. evaluate the overall accuracy at the image level. An expression is a true positive (TP) only if all predictions exactly match the true bounding boxes. Any partial or extra prediction is a false positive (FP). For no-target samples, no prediction is counted as a true negative (TN), otherwise a false positive (FP). Image-level metrics use the same formulations as instance-level metrics but are applied to whole expressions rather than individual objects.
@@ 4 NGDINO
We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges in REC. Our key insight is that explicit numerical reasoning about target counts enhances model performance on these challenges. NGDINO builds upon GDINO [34] while introducing three components: (1) a number prediction head, (2) learnable number-queries with number-guided query selection, and (3) a number cross-attention module.
NGDINO adopts GDINO’s dual-encoder-single-decoder architecture. The image encoder extracts visual features from input images, while the text encoder processes referring expressions. A feature enhancer facilitates cross-modal fusion, and a language-guided query selection module initializes detection queries. The decoder iteratively refines detection queries through self-attention, image cross-attention, and text cross-attention mechanisms, ultimately producing bounding box predictions. We mainly improve the decoder part, as highlighted by the yellow box in Figure 6.
Number prediction head. We introduce a classification module to predict the number of referred objects $N_{pred}$. This prediction head consists of an FFN layer applied to the detection queries $Q_{det}$, followed by mean pooling:
$N_{prob}=\text{softmax}~{}(~{}\text{MeanPool~{}(~{}FFN}~{}(Q_{det}))),$ (3)
$N_{pred}=\text{argmax}~{}(N_{prob}),$ (4)
where $Q_{det}\in\mathbb{R}^{B\times L_{d}\times D}$ with batch size $B$, detection query length $L_{d}$, and feature dimension $D$. Rather than directly predicting arbitrary counts, we quantize the prediction space into five categories $\mathcal{C}=\{0,1,2,3,4+\}$, where $4+$ represents four or more targets. This quantization addresses the long-tailed distribution of target counts in real-world data (which follows Zipfian distribution [2]).
Figure 6: Architecture of a single decoder layer in Number GroundingDINO. Key modifications from GDINO [34] (highlighted in yellow box) include: (1) number prediction head (FFN) to predict target number, (2) number-queries selected through the predicted number, (3) number cross-attention between selected number-queries and detection-queries.
Number-queries and number-guided query selection. We introduce a learnable set of number-queries $Q_{num}\in\mathbb{R}^{B\times L_{n}\times D}$, where $L_{n}$ is the length of number-queries. $Q_{num}$ is randomly initialized and trained to capture numerical patterns. Based on the predicted target number $N_{pred}$, we select a subset of number-queries $Q_{num}^{sel}$:
@@
where $L_{s}$ is the length of selected number-queries. This creates a direct mapping between the predicted number of objects and the specialized queries that encode patterns relevant to that cardinality.
Number cross-attention module. We integrate the number information into the object detection process through a number cross-attention mechanism that operates in parallel with the self-attention in the decoder. In number cross-attention, detection queries $Q_{det}$ serve as queries (Q), while the selected number-queries $Q_{num}^{sel}$ serve as both keys (K) and values (V). The output of this module is added to the self-attention features before subsequent processing in the decoder layer.
Training objective. We adopt GDINO’s bounding box supervision and add a Cross-Entropy loss for number prediction. Ablation studies determine the optimal parameters: selected number-query length $L_{s}=10$ and total number-query length $L_{n}=50$, matching our 5-category quantization scheme.
@@ 5 Experiments
We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs. The specialist models include MDETR [20], GLIP [25], and GDINO [34]. For the LMMs, we evaluate Shikra [7], ONE-PEACE [47], SPHINX-v2 [27], MiniGPT-v2 [5], Ferret [53], Kosmos-2 [38], Griffon [57], Qwen-VL [3], CogVLM [50], and LLaVA [30]. Detailed model specifications and implementation details are provided in the Appendix.
@@
Table 2: Experimental results of fine-tuning baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks. GDINO-T and GDINO-B denote GroundingDINO [34] with Swin-Tiny and Swin-Base backbones. The best results in each group are denoted with bold.
@@
Methods F1inst. Accinst. F1img. Accimg. MDETR [20] 32.60 19.49 19.17 10.81 GLIP [25] 24.23 14.86 16.92 13.29 GDINO-T [34] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [34] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [5] 4.97 2.74 13.56 8.97 LLaVA-v1.5 [30] 6.00 3.63 14.43 11.57 Qwen-VL [3] 14.14 7.61 20.10 11.17
Table 3: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
Methods F1inst. Accinst. F1img. Accimg. MDETR [20] 8.42 4.41 2.99 1.63 GLIP [25] 5.46 3.84 9.20 8.54 GDINO-T [34] 1.18 1.84 3.94 6.35 GDINO-B [34] 1.97 2.23 6.43 7.58 Shikra [7] 0.80 0.52 2.26 1.60 ONE-PEACE [47] 1.02 0.51 2.64 1.34 SPHINX-v2 [27] 1.59 0.80 4.61 2.36 MiniGPT-v2 [5] 2.69 1.36 6.38 3.29 Ferret [53] 3.18 1.62 8.48 4.43 Kosmos-2 [38] 8.06 4.20 8.64 4.52 Griffon [57] 9.16 4.81 16.77 9.18 Qwen-VL [3] 10.91 5.77 18.32 10.08 CogVLM [50] 15.38 8.33 30.73 18.15
5.1 Experimental results
Fine-tuning results. Table 3 illustrates fine-tuning performance across multiple methods on our RefDrone dataset. The consistently poor performance across all approaches highlights the inherent challenges of the proposed dataset. The specialist model MDETR [20] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [34], but struggles with image-level understanding. Conversely, LMMs like Qwen-VL [3] achieve superior image-level comprehension while struggling with instance-level tasks. This performance disparity can be attributed to LMMs’ effective global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints (detailed results by object scale are provided in Appendix Table 9).
Our proposed NGDINO consistently outperforms the baseline GDINO [34] across both backbone architectures. Specifically, NGDINO-B achieves state-of-the-art performance, surpassing GDINO-B by significant margins in instance-level metrics (2.48% improvement in F1inst. and 1.77% in Accinst.) as well as image-level metrics (2.32% in F1img. and 1.63% in Accimg.).
Zero-shot results. Table 3 presents the zero-shot evaluation results, assessing the models’ domain generalization capability. CogVLM [50] demonstrates state-of-the-art performance across metrics. However, several advanced models, including Shikra [7], ONE-PEACE [47], and SPHINX-v2 [27], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy. This restriction significantly impacts their performance in multi-target scenarios.
Additional benchmark evaluations. As presented in Table 5, we evaluate our method on the gRefCOCO [28] benchmark with multi-target and no-target samples. The metric Pr@0.5 measures the percentage of predictions with $F1$ = 1 at IoU $\geq$ 0.5, while N-acc denotes the no-target accuracy. NGDINO-T demonstrates a notable improvement over GDINO-T, especially in the N-acc metric, achieving gains of 4.15% on test A and 1.39% on test B, which assess performance on no-target samples. In contrast, the performance gains in Pr@0.5 (0.36% and 0.74% on test A and B, respectively) are more modest. This can be partially attributed to the relatively simple structure of the multi-target queries in gRefCOCO, where expressions predominantly refer to only one or two objects. MDETR [20] achieves higher Pr@0.5 on the test A set. However, this is likely because MDETR tends to produce more outputs, resulting in lower N-acc performance due to an increased number of false positives on no-target examples.
Table 5 demonstrates our method’s effectiveness on the aerial view RSVG [56] benchmark with multi-target samples, where NGDINO-T consistently outperforms previous SOTA methods.
@@ 5.1 Experimental results
Table 4: Experimental results on gRefCOCO [28] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
@@
Methods testA testB Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$ MDETR* [20] 50.0 34.5 36.5 31.0 UNINEXT* [52] 46.4 49.3 42.9 48.2 GDINO-T [34] 45.69 79.02 44.83 76.69 NGDINO-T 46.05 83.17 45.57 78.08

### v3

@@ 3.5 Evaluation metrics
To properly assess performance on multi-target expressions, we introduce instance-level metrics alongside traditional image-level ones. Standard image-level metrics, which treat the entire set of predictions for an expression as a single entity, are insufficient for our task. They cannot granularly penalize a model for missing individual objects within a large group. Our proposed instance-level metrics address this limitation by providing a fine-grained assessment of a model’s ability to localize each individual target.
Instance-level metrics Accinst. and F1inst.: These metrics evaluate performance at the individual object level. We match each predicted bounding box to a ground-truth (GT) box. A prediction with an IoU $\geq$ 0.5 with a GT box is a true positive (TP). Unmatched predictions are false positives (FP), and unmatched GT boxes are false negatives (FN). For “no-target” samples, the absence of any prediction is a true negative (TN), while any prediction is an FP. Accuracy and F1-score are then computed as:
@@
$\text{F1}_{inst.}=\frac{2\cdot\text{TP}}{2\cdot\text{TP}+\text{FP}+\text{FN}}.$ (2)
Image-level metrics Accimg. and F1img.: These metrics assess performance on the entire expression, enforcing a stricter success criterion. For a given expression, a prediction is a true positive (TP) only if the set of predicted boxes perfectly matches the set of ground-truth boxes. Any mismatch (e.g., missing, extra, or inaccurate boxes) results in a false positive (FP). TN and FP for “no-target” samples are defined as above. The metrics are calculated using the same formulas but on an image-wide basis.
@@ 4 NGDINO
We introduce Number GroundingDINO (NGDINO), a novel architecture designed to address the challenges of multi-target and no-target referring expression comprehension (REC). Our core insight is that explicit numerical reasoning about target counts is crucial for accurately grounding such expressions. NGDINO builds upon the strong foundation of GDINO [gdino], inheriting its dual-encoder, single-decoder structure. Our contributions are concentrated within the decoder, as highlighted in Figure 6, where we introduce three key components: (1) a number prediction head, (2) learnable number-queries with number-guided query selection, and (3) a number cross-attention module.
Number prediction head. To explicitly reason about the number of targets, we append a lightweight number prediction head to the decoder. This head takes the detection queries $Q_{det}$ as input, processes them through a Feed-Forward Network (FFN), and applies mean pooling to produce a probability distribution over target counts:
$N_{prob}=\text{softmax}\left(\text{MeanPool}\left(\text{FFN}(Q_{det})\right)\right),$ (3)
$N_{pred}=\text{argmax}(N_{prob}),$ (4)
where $Q_{det}\in\mathbb{R}^{B\times L_{d}\times D}$ are the detection queries for a batch size $B$, query length $L_{d}$, and feature dimension $D$. To manage the long-tailed, Zipfian-like distribution of target counts [axtell2001zipf], we discretize the output space into five bins: $\mathcal{C}=\{0,1,2,3,4+\}$, where $4+$ represents four or more targets. This approach transforms the open-ended regression problem into a stable multi-class classification task.
Number-guided query selection. To inject numerical priors into the decoding process, we introduce learnable number-queries, $Q_{num}\in\mathbb{R}^{B\times L_{n}\times D}$, where $L_{n}$ is the length of number-queries. These queries are trained to encode distinct numerical patterns. The predicted number, $N_{pred}$, then acts as an index to select a dedicated slice of these queries:
@@
where $L_{s}$ is the length of number-queries per category. This mechanism creates a mapping between the predicted cardinality and queries encoding patterns.
Number cross-attention module. The selected number-queries, $Q_{num}^{sel}$, are fused with the detection queries, $Q_{det}$, via a number cross-attention module. This module operates in parallel with the self-attention layer within the decoder. The detection queries serve as the query (Q), while the selected number-queries serve as both key (K) and value (V). The output of this module is added to the output of the self-attention layer, enriching the detection queries with numerical context before they are passed to subsequent layers.
Training objective. The model is trained end-to-end by augmenting the GDINO losses (for bounding box regression and label classification) with a Cross-Entropy loss for the number prediction task. Hyperparameters, including the selected number-query length ($L_{s}=10$) and total number-query length ($L_{n}=50$, representing 5 categories $\times$ 10 queries), were determined via ablation studies in the Appendix.
@@ 5 Experiments
We establish a comprehensive benchmark comprising 26 representative methods capable of performing REC tasks: 3 specialized REC methods, 7 task-specific LVLMs for REC, 12 general LVLMs with varying parameter scales, and 4 closed-source API models. To evaluate our proposed NGDINO, we conduct experiments on our RefDrone dataset and two public benchmarks: gRefCOCO [grefcoco] and RSVG [zhan2023rsvg].
@@
Table 2: Experimental results of zero-shot baselines on RefDrone benchmark. The best results in each group are denoted with bold.
Categories Methods Params Time F1inst. Accinst. F1img. Accimg. Specialized REC Methods $\text{MDETR}_{\text{ResNet101}}$ [mdetr] 0.19B 2021/04/26 8.42 4.41 2.99 1.63 $\text{GLIP}_{\text{Swin-Tiny}}$ [glip] 0.15B 2021/12/06 5.46 3.84 9.20 8.54 $\text{GroundingDINO}_{\text{Swin-Tiny}}$ [gdino] 0.17B 2023/05/09 1.18 1.84 3.94 6.35 $\text{GroundingDINO}_{\text{Swin-Base}}$ [gdino] 0.23B 2023/05/09 1.97 2.23 6.43 7.58 Specialized LVLMs for REC Shikra [chen2023shikra] 7B 2023/07/03 0.80 0.52 2.26 1.60 $\text{ONE-PEACE}_{\text{Grounding}}$  [one_peace] 4B 2023/07/20 1.02 0.51 2.64 1.34 Kosmos-2 [peng2023kosmos] 1.6B 2023/10/30 8.06 4.20 8.64 4.52 $\text{CogVLM}_{\text{Grounding}}$ [wang2023cogvlm] 7B 2023/11/20 15.38 8.33 30.73 18.15 Griffon [zhan2025griffon] 13B 2023/12/06 9.16 4.81 16.77 9.18 Ferret [you2023ferret] 7B 2023/12/14 3.18 1.62 8.48 4.43 Rex-Omni [jiang2025detect] 3B 2025/10/15 54.06 37.10 43.90 28.52 General LVLMs 0 ~ 7B $\text{DeepSeek-VL2}_{\text{Tiny}}$ [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 3B 2024/12/13 2.35 1.84 4.51 4.98 Qwen2.5-VL [bai2025qwen2] 3B 2025/02/20 40.00 25.06 38.22 23.89 Qwen3-VL [qwen3technicalreport] 4B 2025/10/15 48.68 32.31 47.78 32.28 General LVLMs 7B ~ 10B Qwen-VL [bai2023qwen] 7B 2023/08/12 10.91 5.77 18.32 10.08 MiniGPT-v2 [chen2023minigpt] 7B 2023/10/13 2.69 1.36 6.38 3.29 InternVL2.5 [internvl2_5] 8B 2024/12/05 0.58 1.14 1.79 4.2 Qwen2.5-VL [bai2025qwen2] 7B 2025/02/20 42.68 27.20 41.46 26.52 GLM 4.1V [vteam2025glm45vglm41vthinkingversatilemultimodal] 9B 2025/07/01 24.39 14.05 30.57 18.74 Ovis2.5 [lu2025ovis2] 9B 2025/08/19 4.24 2.20 14.56 8.39 $\text{MiMo-VL}_{\text{RL}}$ [coreteam2025mimovltechnicalreport] 7B 2025/08/21 19.75 11.22 30.24 18.79 InternVL3.5 [wang2025internvl3_5] 8B 2025/08/26 8.17 4.48 19.73 11.8 Qwen3-VL [qwen3technicalreport] 8B 2025/10/15 51.66 35.13 46.36 31.85 General LVLMs 10B ~ 30B SPHINX-v2 [lin2023sphinx] 13B 2023/11/17 1.59 0.80 4.61 2.36 $\text{DeepSeek-VL2}_{\text{Small}}$ [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 16B 2024/12/13 34.95 21.34 36.15 22.81 DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 27B 2024/12/13 25.31 14.88 29.46 19.08 Qwen3-VL [qwen3technicalreport] 30B 2025/10/15 56.88 39.98 51.09 35.58 General LVLMs > 30B GLM 4.5V [vteam2025glm45vglm41vthinkingversatilemultimodal] 106B 2025/08/11 35.95 21.97 40.21 25.41 Qwen3-VL [qwen3technicalreport] 235B 2025/10/04 58.79 41.93 52.16 36.89 Closed-Source (APIs) Models DINO-XSeek [ren2024dino] 2025/03/11 54.47 37.46 46.11 30.14 Gemini 2.5 Pro [comanici2025gemini] 2025/03/25 3.45 1.91 8.11 4.90 Seed1.5-VL [guo2025seed1] 2025/05/12 43.52 27.84 37.27 23.03 Qwen3-VL-Plus [qwen3technicalreport] 2025/09/22 58.11 41.32 50.98 36.18 $\dagger$} State-of-the-art method.   $\ddagger$} Models only predict a single bounding box, limiting multi-target performance. * GPT and Claude are excluded due to output bounding-box format failures.
5.1 Zero-shot results.
As presented in Table 2, we evaluate all models in a zero-shot setting to assess their generalization capabilities to the drone scenes.
Overall performance. Among all models, the open-source Qwen3-VL (235B) [qwen3technicalreport] achieves state-of-the-art performance, with 58.79% F1inst., 41.93% Accinst., 52.16% F1img. , and 36.89% Accimg.. These results surpass not only other open-source models but also the specialized closed-source API, DINO-XSeek [ren2024dino] (54.47% F1inst.), which is explicitly designed for REC tasks.
Analysis of specialized methods. Specialized REC models (e.g., MDETR [mdetr], GLIP [glip], GroundingDINO [gdino]) show limited zero-shot transfer to RefDrone, with F1inst. generally below 10%. This likely stems from pre-training on general-domain datasets, which restricts their cross-domain adaptability. Task-specific LVLMs for REC show varied performance. Models limited to predicting a single bounding box, like Shikra [chen2023shikra], ONE-PEACE [one_peace], and SPHINX-v2 [lin2023sphinx], fail in multi-target scenarios, resulting in F1inst. scores below 2%. In contrast, recent Rex-Omni [jiang2025detect] achieves competitive performance (54.06% F1inst. / 43.90% F1img.), approaching the state-of-the-art.
Analysis of general LVLMs. A strong correlation is observed between model scale and zero-shot REC performance in general-purpose LVLMs. For instance, the Qwen3-VL series shows consistent improvement with scale: the 4B model achieves 48.68% F1inst., which increases to 51.66% for the 8B model and 56.88% for the 30B model. This scaling trend suggests that larger model capacity directly enhances localization abilities without task-specific tuning. A notable exception is DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels], where the small variant achieves better performance.
Performance of closed-source APIs. Commercial models do not consistently outperform open-source alternatives. DINO-XSeek [ren2024dino] achieves 54.47% F1inst., lagging behind Qwen3-VL-235B. Other prominent APIs, including Gemini 2.5 Pro, yield low scores (F1inst. < 4%). Furthermore, models from the GPT and Claude families failed to produce outputs in the required bounding-box format. This disparity indicates that many leading commercial models are not optimized for fine-grained localization tasks like REC.
Table 3: Results of fine-tuning baselines on RefDrone benchmark.
@@
Methods F1inst. Accinst. F1img. Accimg. GLIP-T [glip] 56.92 40.39 41.31 28.88 GDINO-T [gdino] 67.64 51.55 54.54 39.63 NGDINO-T (Ours) 71.11 55.52 56.51 41.20 GDINO-B [gdino] 69.75 53.95 56.95 41.81 NGDINO-B (Ours) 72.51 57.22 57.84 42.54
5.2 Fine-tuning results.
Performance on the RefDrone Dataset. Table 3 presents the fine-tuning performance across specialized REC methods for 50 epochs on RefDrone dataset. Our proposed NGDINO exhibits consistent gains over the GDINO baseline. Specifically, the Swin-Tiny models NGDINO-T surpasses GDINO-T by 3.47%, 3.97%, 1.97%, and 1.57% in F1inst., Accinst., F1img., and Accimg.. Our NGDINO-B achieves the highest performance: 72.51% F1inst., 57.22% Accinst., 57.84% F1img. , and 42.54% Accimg., outperforming the state-of-the-art LVLM Qwen3-VL-235B by a significant margin. Beyond the performance advantage, NGDINO requires orders of magnitude fewer parameters than LVLMs, making it suitable for deployment on resource-constrained edge platforms such as drones.
@@
Table 4: Experimental results on gRefCOCO [grefcoco] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
@@
Methods testA testB Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$ MDETR* [mdetr] 50.0 34.5 36.5 31.0 UNINEXT* [yan2023universal] 46.4 49.3 42.9 48.2 GDINO-T [gdino] 45.69 79.02 44.83 76.69 NGDINO-T 46.05 83.17 45.57 78.08
Performance on public benchmarks. To further validate the effectiveness of our improvements over GDINO in handling multi-target scenarios, we evaluate NGDINO on two benchmarks that contain multi-target samples: gRefCOCO [grefcoco] and RSVG [zhan2023rsvg]. Models are fine-tuned for 5 epochs on both benchmarks. Table 4 presents results on gRefCOCO [grefcoco], which includes both multi-target and no-target samples. We adopt two metrics: Pr@0.5, measuring the percentage of predictions achieving $F1$ = 1 at IoU $\geq$ 0.5, and N-acc, denoting accuracy on no-target samples. NGDINO-T demonstrates improvements over GDINO-T in no-target detection, with N-acc gains of 4.15% and 1.39% on test A and test B, respectively. The improvements in Pr@0.5 are more modest (0.36% and 0.74% on test A and B), which can be attributed to the relatively simple multi-target structure in gRefCOCO, where expressions predominantly reference only one or two objects. While MDETR [mdetr] achieves higher Pr@0.5 on test A, this comes at the cost of lower N-acc due to its tendency to produce excessive outputs, leading to false positives on no-target samples. On the more complex aerial-view RSVG [zhan2023rsvg] benchmark (Table 5), NGDINO-T consistently outperforms prior state-of-the-art methods across all metrics. These results confirm that our proposed modifications effectively enhance performance on diverse multi-target REC tasks.
5.3 Ablation studies
Analysis of NGDINO components. Table 6 presents the analysis of each component in NGDINO. With only the number prediction head, the model achieves improvements, improving F1inst. from 67.64% to 69.73% (+2.09%). This demonstrates that the auxiliary number prediction task enhances the model’s grounding capability. The number cross-attention mechanism alone yields more modest gains, improving F1inst. to 68.88% (+1.24%). This improvement can be partially attributed to the additional parameters introduced in the decoder. The most significant performance boost occurs when combining both components, where F1inst. reaches 71.11% (+3.47%). This substantial accuracy improvement comes at a negligible cost to efficiency, with only a minor decrease in inference speed (from 13.5 to 12.3 FPS).
Number prediction accuracy. We evaluate the number prediction head’s performance. The head achieves a mean absolute error (MAE) of 0.21, indicating that its numerical predictions are highly precise and closely align with the ground truth counts. This strong performance in number prediction is a key factor in the model’s ability to handle multi-target expressions. The overall accuracy for predicting the correct number of instances is 75.3%.